OpenStudy (anonymous):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2

OpenStudy (anonymous):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? The equation of a line is given by: y=mx+c m = gradient or slope == change in y/change in x c = y_intercept (where the line crosses x = 0) First calcualte the gradient m= (12--2)/(-5-2) = 14/-7= -2 Step 2: Now find the y-intercept: We use the point (−5, 12) 12 = (-2*−5)+b 12 = 10+b b= 12-10 = 2 equation is for the line is: y= -2x+2

OpenStudy (anonymous):

so is it A? @twicker